OpenStudy (anonymous):

Find a unit vector orthogonal to −5 I + 2 J + 10 K and 0 I + 1 J + 0 K: ??? I + ??? J + ??? K

OpenStudy (anonymous):

Use cross product.

OpenStudy (anonymous):

didnt work

OpenStudy (anonymous):

sure it does, [-10, 0, -5] is the answer.

OpenStudy (anonymous):

i j k -5 2 10 0 1 0 =-10i+0j-5k which is what YTheManifold stated

OpenStudy (anonymous):

i did it but only the 0j is right

OpenStudy (anonymous):

check the answer by calculating the scalar product to see if it is orthogonal

OpenStudy (anonymous):

what difference does that makes !?

OpenStudy (anonymous):

or you could DOT them which will be=0 a*b/||a||||b||=cos(x) 0=cos(x) cos^-1(0)=x x=pi/2 or (-3pi/2) which is orthogonal and if the cos^-1(x)=1 or -1 then they are parallel

OpenStudy (anonymous):

its not parallel i know that too !

OpenStudy (anonymous):

plus i don't need those properties

OpenStudy (anonymous):

the unit vector is -10i+0j-5k/sqrt(100+25) thats a unit vector does that look better? 5sqrt(5) -2/sqrt(5)i+0j+-1/sqrt(5)k

OpenStudy (anonymous):

yeah it does haha thanX

OpenStudy (anonymous):

sweet, welcome